OpenStudy (anonymous):

can you help me find the derivative of f(x)-(x^2+2)^3(x^3+3)^2

OpenStudy (anonymous):

\[f(x)= (x^2+2)^3(x^3+2)^2\]

OpenStudy (anonymous):

http://www.mathway.com/

OpenStudy (anonymous):

f'(x) = \[36x ^{3}(x ^{2}+2)^{2}(x ^{3}+2)\]

OpenStudy (anonymous):

sorry... i m giving wrong answer... i think you can try using math lab

OpenStudy (anonymous):

I would use the product rule d/dx[f(g(x))] = f'(x)g(x)+g'(x)f(x), letting \[f(x)=(x^2+2)^3\] and \[g(x)=(x^3+2)^2\] Use the chain rule to solve for f' and g' d/dx[u^n]=n(u^n-1)u'